OpenStudy (anonymous):

Consider the function y = -2-3cos(x+pi). What effect does the -2 have on the original graph?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

vertical shift down two units horizontal shift right two units horizontal shift left two units vertical stretch by factor of 2

OpenStudy (anonymous):

I think its B?

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

i think ur answer is correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@AntarAzri

OpenStudy (anonymous):

anyone else think its B?

OpenStudy (anonymous):

y-Intercept ?

OpenStudy (anonymous):

it has to be one of the options

OpenStudy (anonymous):

\(cos(x+1\pi)=0\) ?

OpenStudy (anonymous):

sorry about the : \(1\pi\)

OpenStudy (anonymous):

its ok umm I don't know what you mean with the equation I haven't gotten into that yet or well at least they haven't taught me that

OpenStudy (anonymous):

vertical shift down two units horizontal shift right two units horizontal shift left two units vertical stretch by factor of 2

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

Let me investigate

OpenStudy (anonymous):

ok btw I'm taking a timed practice set of problems and have three minutes for this question so..

OpenStudy (anonymous):

Btw, I'm not in exam and i have a lot of time to think about science

OpenStudy (anonymous):

Might help you : http://www.wolframalpha.com/input/?i=Plot%28y+%3D+-2-3cos%28x%2Bpi%29%29

OpenStudy (anonymous):

okayyy so it would be A?

OpenStudy (anonymous):

@AntarAzri

OpenStudy (anonymous):

time moves fast in the real world btw

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

anyone know trig?

OpenStudy (wolf1728):

I made some graphs of the equation They are in this order -2 equation unchanged -3 in place of the -2 -1 in place of the -2

OpenStudy (anonymous):

so vertical shift down 2 units?

OpenStudy (wolf1728):

I'm looking at these right now -2 the maximum is 1 min -5 -1 the maximum is 4 min -2 -3 the maximum is 6 min 0

OpenStudy (anonymous):

ok well while you look I have one last question for the day "Which function's graph has asymptotes at values x = +/- npi? 1. y = csc x 2. y = tan x 3. y = cos x 4. y = cot x 1 only 2 only 1 and 3 only 1 and 4 only I think its either 1 or 1 and 3 only

OpenStudy (anonymous):

@AntarAzri

OpenStudy (wolf1728):

GEEZ - I have graphs of all 6 trig functions but it asks for asymptotes at +/- PI?

OpenStudy (anonymous):

+/- NPi

OpenStudy (wolf1728):

and what is NPi ??

OpenStudy (anonymous):

integer times pi on x axis

OpenStudy (wolf1728):

GEEZ can't they just ask for degrees ??????????????

OpenStudy (anonymous):

I think like 2pi

OpenStudy (wolf1728):

So they want it in radians? They could have said that!!

OpenStudy (anonymous):

lol I know

OpenStudy (anonymous):

so what do u think? 1 only? 1 and 4 maybe?

OpenStudy (wolf1728):

I'll post the max and min of each OR here's a link to the graphs: http://1728.org/trigtutr.htm

OpenStudy (wolf1728):

You'll have to scroll to the bottom.

OpenStudy (anonymous):

ok umm they all deal with degrees and numbers I'm looking for the ones with pi

OpenStudy (wolf1728):

So would 1 PI be 57.2957795131 ??

OpenStudy (anonymous):

I guess

OpenStudy (wolf1728):

It appears that the asymptotes appear at 0° 90° 180° 270° and 360° I don't think any of those have a neat integer value of PI do they?